OpenStudy (anonymous):

One number is 8 more than another. The sum of the two numbers is 26. What are the numbers?

OpenStudy (anonymous):

I will fan and medal!

Vocaloid (vocaloid):

Hint: N = the first number N + 8 = the second number N + N + 8 = 26 Solve for N

OpenStudy (anonymous):

2N+8=26 2N=18 N=9

Vocaloid (vocaloid):

Good, now find N + 8

OpenStudy (anonymous):

17

Vocaloid (vocaloid):

Good do your two numbers are 17 and 9

OpenStudy (anonymous):

Can you help me with 2 more?

Vocaloid (vocaloid):

I guess yea

OpenStudy (anonymous):

\[\frac{ 5 }{ 6 }*\frac{ 1 }{ 2 }*\frac{ 3 }{ 25 }=\]

OpenStudy (anonymous):

The answer has to be simplified.

OpenStudy (anonymous):

\[=\frac{ 15 }{ 300 }\]

Vocaloid (vocaloid):

The 25 can be changed to 5 and the 5 can be changed to a 1

Vocaloid (vocaloid):

Oh, well, 15/300 = 1/20

OpenStudy (anonymous):

1/20 was right

Vocaloid (vocaloid):

Right, all I did was divide the numerator and denominator by 15

OpenStudy (anonymous):

Right! Last one!\[\frac{ 30x }{ 37 }*\frac{ 23 }{ 10}\]

OpenStudy (anonymous):

It comes out to 690/370

Vocaloid (vocaloid):

The 30 and the 10 cancel to get 3 and 1

OpenStudy (anonymous):

69/37

Vocaloid (vocaloid):

Almost, don't forget the x

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

YES!!!! THANK YOU SO MUCH!!!!!!

Vocaloid (vocaloid):

69x/37

Vocaloid (vocaloid):

Great!!!

OpenStudy (anonymous):

Great....I have 2 more....can you help?

Vocaloid (vocaloid):

K

OpenStudy (anonymous):

\[\left(-\begin{matrix}5 \\ 6\end{matrix}\right)^{2}\]

Vocaloid (vocaloid):

Square the numerator and denominator

OpenStudy (anonymous):

- 25/36

Vocaloid (vocaloid):

The negative sign cancels out to get a positive answer

Vocaloid (vocaloid):

25/36

OpenStudy (anonymous):

okay

Vocaloid (vocaloid):

Remember, (-1) squared is 1

OpenStudy (anonymous):

Last One! I promise! (7x+9)-(-2x^2-5x+9)

Vocaloid (vocaloid):

Distribute the negative sign first

OpenStudy (anonymous):

I got it right! Thank you so much for your help!

Vocaloid (vocaloid):

Awesome

OpenStudy (anonymous):

Your great!